# Degree-one foliations on complete intersections

### v1

@@ 2.3. Degree-zero distributions on the projective space
In [1], the authors classify distributions of codimension one and degree zero on ${\mathbb{P}^{n}}$. There is a stratification according to the class of the $1$-form defining the distribution. The argument boils down to the classification of skew-symmetric matrices and the nesting of secant varieties to the Grassmannian of lines in ${\mathbb{P}}(\textstyle\bigwedge^{2}{\mathbb{C}}^{n+1})$. In higher codimension, the situation is surprisingly simpler.
@@
Let $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(q+1))$ a $q$-form defining $\mathcal{D}$. By [11, Theorem A], there exists a choice of homogeneous coordinates such that either $\omega=\alpha\wedge dx_{0}\wedge\cdots\wedge dx_{q-2}$ for some $\alpha$ a $1$-form with linear coefficients, or $\omega=i_{v}dx_{0}\wedge\cdots\wedge dx_{q}$, for some linear vector field $v$. Imposing $i_{\rm rad}\omega=0$, for ${\rm rad}=\sum_{j}x_{j}\frac{\partial}{\partial x_{j}}$ the radial vector field, we must have $\omega=i_{\rm rad}dx_{0}\wedge\cdots\wedge dx_{q}$, up to a scalar multiple. Therefore, $\mathcal{D}$ is integrable. ∎
@@ 3. Extensions and restrictions of foliations on complete intersections
Let $U_{k}$ be the open subset of twisted $q$-forms in $H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(k))$ whose singular set has codimension greater than or equal to two and let $X\subset{\mathbb{P}^{n}}$ be a smooth complete intersection.
@@
We say that $\omega\in U_{k}$ is transversal in codimension one to $X$ if $i^{*}\omega$ has a singular set of codimension greater than or equal to two, where $i\colon X\rightarrow{\mathbb{P}^{n}}$ is the natural inclusion. A singular holomorphic foliation $\mathcal{F}$ of codimension $q$ is transversal in codimension one to $X$, if a twisted $q$-form that determines it is transversal in codimension-one to $X$.
If a foliation $\mathcal{F}$ is transversal in codimension one to $X$, the restriction to $X$ of its twisted $q$-form naturally satisfies the decomposability and integrability hypotheses and, therefore, defines a foliation $\mathcal{G}$ on $X$.
@@ 3.1. Codimension-one foliations
In general, for logarithmic forms as in (2.1), we have the commutative diagram
[…]
which projects all logarithmic forms of ${\mathbb{P}^{n}}$ onto $X$ surjectively, where $i\colon X\rightarrow{\mathbb{P}^{n}}$ is the natural inclusion map. Therefore, all foliations generated by logarithmic forms extend.
@@
As seen above, if $\mathcal{F}$ is generated by a logarithmic form, it has an extension. Furthermore, the logarithmic and rational components can be determined by the closure of the image of a rational map as (3.1).
Suppose $\mathcal{F}$ is in a logarithmic component $\textsf{Log}(X,d_{1},d_{2},\dots,d_{k})\subset{\sf Fol}(X,\sum d_{i}-2)$ and is an adherence value of the map $\Phi$. In this case, there exists a sequence of foliations $\{\mathcal{F}_{i}\}_{i=0}^{\infty}$ in the logarithmic component such that $\mathcal{F}_{i}\rightarrow\mathcal{F}$ and each $\mathcal{F}_{i}$ is defined by a logarithmic form. Since each foliation $\mathcal{F}_{i}$ extends, there exists a sequence of foliations $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ in the logarithmic component $\textsf{Log}({\mathbb{P}^{n}},d_{1},d_{2},\dots,d_{k})$, where ${\left.\kern-1.2pt\mathcal{G}_{i}\vphantom{\big{|}}\right|_{X}}=\mathcal{F}_{i}$ and each $\mathcal{G}_{i}$ is determined by a logarithmic form $\omega_{i}$. Since the integrability condition determines a proper closed set in the class space $\mathbb{P}H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{1}(k)),$ the set of points adhering to the sequence $([\omega_{i}])_{i=0}^{\infty}$ is nonempty and each point in this set is integrable. Thus, given $[\omega]$ as one of these adherent points, its restriction to $X$ determines $\mathcal{F}$ by definition. Furthermore, we have that $\operatorname{codim}(\omega)\geq 2$, otherwise, $\operatorname{codim}(\operatorname{rest}(\omega))\leq 1$ and $\operatorname{rest}(\omega)$ would not determine $\mathcal{F}$. Therefore, $\omega$ determines a logarithmic foliation $\mathcal{G}$ on ${\mathbb{P}^{n}}$ whose restriction is $\mathcal{F}$. ∎
@@ 3.2. Codimension two foliations
For codimension-two foliations of degree zero, we have the following results regarding extensions.
@@
Let $\mathcal{F}$ be a codimension-two foliation on $X$ of degree zero. In particular, $\mathcal{F}$ is a codimension-two distribution on $X$ of degree zero. Applying [14, Corolário 2.3.3], we see that $\mathcal{F}$ has a unique extension as a codimension-two distribution to ${\mathbb{P}^{n+s}}$ of degree zero. Finally, by Proposition 2.4, we conclude that the extension of $\mathcal{F}$ is integrable. ∎
@@
It follows by Theorem 3.2, if $\deg(X)\geq 3$ and, for quadrics, by [3, Proposition 3.18]. ∎
@@ 4.1. Tangential foliations
A smooth projective variety $X$ is called uniruled if, for every general point $x\in X$, there exists a non-constant morphism $f\colon{\mathbb{P}^{1}}\to X$ such that the rational curve $C=f({\mathbb{P}^{1}})$ passes through $x$ and $f^{*}{TX}$ is generated by global sections; such morphisms are called free. The scheme $\operatorname{Mor}({\mathbb{P}^{1}},X)$ parametrizing morphisms from ${\mathbb{P}^{1}}$ to $X$ is smooth around a free morphism $f$ and its tangent space is naturally isomorphic to $H^{0}(X,f^{*}{TX})$. We refer to [12, 18] for further details. A particular result we will need is the following.
@@
Let $f\colon{\mathbb{P}^{1}}\to X$ be a free morphism and $Z\subset X$ be a closed subvariety of codimension at least two. Then, $f^{\prime}({\mathbb{P}^{1}})\cap Z=\emptyset$ for a general deformation $f^{\prime}$ of $f$.
@@
This is a particular case of [18, II, Proposition 3.7]. ∎
@@
See [12, Proposition 4.14] for proof of the existence of $X^{\rm free}$ over any algebraically closed field of characteristic zero. If the base field is uncountable, then $X^{\rm free}\neq\emptyset$. ∎
@@
and the natural projection $\pi:M\times{\mathbb{P}^{1}}\longrightarrow M$. We have an isomorphism $T_{M}\cong\pi_{*}\mathrm{ev}^{*}T_{X}$, and so the inclusion $T_{\mathcal{F}}\subset T_{X}$ induces a morphism $i:\pi_{*}\mathrm{ev}^{*}T_{\mathcal{F}}\hookrightarrow T_{M}$. We define the tangential foliation of $\mathcal{F}$ on $M$ by the saturation of the image of this morphism $i$ in $T_{M}$. In [21], the authors show how the geometry of $\mathcal{F}_{\rm tang}$ impacts the geometry of $\mathcal{F}$. We recall below two of their results.
@@
[21, Proposition 4.10] Let $X$ be a uniruled smooth projective variety and let $\mathcal{F}$ be a codimension one foliation on $X$. If all leaves of $\mathcal{F}_{\rm tang}$ are algebraic, then there exists an algebraically integrable foliation $\mathcal{G}$ contained in $\mathcal{F}$ such that $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})=H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})$ for any sufficiently general $f\in M$. In particular, $\mathcal{F}$ is the pullback under a rational map of a foliation on a lower-dimensional manifold.
@@
[21, Theorem C] Suppose that $X$ is simply connected, $\operatorname{codim}\mathcal{F}=1$, the general leaf of $\mathcal{F}_{\rm tang}$ is not algebraic, and a general $f\in M$ nontrivially and transversely all the algebraic hypersurfaces invariant by $\mathcal{F}$. Then $\mathcal{F}$ is defined by a closed rational $1$-form without divisorial components in its zero set.
@@
Due to Lemma 4.1, we may choose $f^{\prime}$ so that $f^{\prime}({\mathbb{P}^{1}})$ does not intersect $\operatorname{Sing}(\mathcal{F})$. Aiming for a contradiction, suppose $f^{\prime}$ is tangent to $\mathcal{F}$. Recall Bott’s partial connection $\nabla^{B}\colon{T_{\mathcal{F}}}\otimes{N_{\mathcal{F}}}\to{N_{\mathcal{F}}}$ that is defined as follows. Let $v\in{T_{\mathcal{F}}}$ and $\overline{w}\in{N_{\mathcal{F}}}$ be local sections, and $w\in{TX}$ a lift of $w$. Then $\nabla^{B}_{v}(\overline{w})=\overline{[v,w]}$. Since $f^{\prime}$ is tangent to $\mathcal{F}$, we get that $f^{\prime*}\nabla^{B}$ is a holomorphic connection on the vector bundle $f^{\prime*}{N_{\mathcal{F}}}$ on ${\mathbb{P}^{1}}$. By Weil’s Theorem [4, Theorem 10], $f^{\prime*}{N_{\mathcal{F}}}$ must be trivial. Using that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, we conclude that $c_{1}({N_{\mathcal{F}}})=0$, hence $\mathcal{F}$ is given by a global holomorphic $p$-form $\omega\in H^{0}(X,\Omega^{p}_{X})$, $p=\operatorname{codim}\mathcal{F}$. On the other hand, $X$ is uniruled with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$ implies that $X$ is Fano, hence $h^{0}(X,\Omega^{p}_{X})=h^{0,p}(X)=0$, which is a contradiction. ∎
@@
Suppose $X\subset{\mathbb{P}^{N}}$ is a complete intersection given by equations of degrees $d_{1},\dots,d_{s}$. If $d_{1}+\cdots+d_{s}\leq N-1$ then $X$ is covered by lines, see [12, Proposition 2.13]. Then, the previous corollary applies.

### v2

@@ 2.3. Degree-zero distributions on the projective space
In [ACM-dist], the authors classify distributions of codimension one and degree zero on $\mathbb{P}^{n}$}. There is a stratification according to the class of the $1$-form defining the distribution. The argument boils down to the classification of skew-symmetric matrices and the nesting of secant varieties to the Grassmannian of lines in ${\mathbb{P}}(\textstyle\bigwedge^{2}{\mathbb{C}}^{n+1})$. In higher codimension, the situation is surprisingly simpler.
@@
Let $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(q+1))$ be a $q$-form defining $\mathcal{D}$. By [Med, Theorem A], there exists a choice of homogeneous coordinates such that either $\omega=\alpha\wedge dx_{0}\wedge\cdots\wedge dx_{q-2}$ for some 1-form $\alpha$ with linear coefficients, or $\omega=i_{v}dx_{0}\wedge\cdots\wedge dx_{q}$, for some linear vector field $v$. Imposing $i_{\rm rad}\omega=0$, for ${\rm rad}=\sum_{j}x_{j}\frac{\partial}{\partial x_{j}}$ the radial vector field, we must have $\omega=i_{\rm rad}dx_{0}\wedge\cdots\wedge dx_{q}$, up to a scalar multiple. Therefore, $\mathcal{D}$ is integrable. ∎
@@ 3. Extensions and restrictions of foliations on complete intersections
Let $U_{k}$ be the open subset of twisted $q$-forms in $H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{q}(k))$ whose singular set has codimension greater than or equal to two, and let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection.
@@
We say that $\omega\in U_{k}$ is transversal in codimension one to $X$ if $i^{*}\omega$ has a singular set of codimension greater than or equal to two, where $i\colon X\rightarrow{\mathbb{P}^{N}}$ is the natural inclusion. A singular holomorphic foliation $\mathcal{F}$ on $\mathbb{P}^{N}$} of codimension $q$ is transversal in codimension one to $X$, if a twisted $q$-form that determines it is transversal in codimension one to $X$.
If a foliation $\mathcal{F}$ on $\mathbb{P}^{N}$} is transversal in codimension one to $X$, the restriction to $X$ of its twisted $q$-form naturally satisfies the decomposability and integrability hypotheses and, therefore, defines a foliation $\mathcal{G}$ on $X$.
@@ 3.1. Codimension-one foliations
In general, for logarithmic 1-forms as in (2.1), we have the commutative diagram
(3.1) ${\Sigma\times\left(\prod^{k}_{i=1}\mathbb{P}H^{0}({\mathbb{P}^{N}},\mathcal{O}_{{\mathbb{P}^{N}}}(d_{i}))\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\Phi_{{\mathbb{P}^{N}}}}$$\scriptstyle{\text{id}\times\operatorname{rest}_{0}}$${\mathbb{P}H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{1}(\sum d_{i}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{*}}$${\Sigma\times\left(\prod^{k}_{i=1}\mathbb{P}H^{0}(X,{\mathcal{O}}_{X}(d_{i}))\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\Phi_{X}}$${\mathbb{P}H^{0}(X,\Omega_{X}^{1}(\sum d_{i}))}$
which projects all logarithmic 1-forms of $\mathbb{P}^{N}$} onto $X$ surjectively, where $i\colon X\rightarrow{\mathbb{P}^{N}}$ is the natural inclusion map. Therefore, all foliations generated by logarithmic 1-forms extend.
@@
As seen above, if $\mathcal{F}$ is generated by a logarithmic 1-form, it has an extension. Furthermore, the logarithmic and rational components can be determined by the closure of the image of a rational map as (3.1).
Suppose $\mathcal{F}$ is in a logarithmic component $\textsf{Log}(X,d_{1},d_{2},\dots,d_{k})\subset{\sf Fol}(X,\sum d_{i}-2)$ and is an adherence value of the map $\Phi$. In this case, there exists a sequence of foliations $\{\mathcal{F}_{i}\}_{i=0}^{\infty}$ in the logarithmic component such that $\mathcal{F}_{i}\rightarrow\mathcal{F}$ and each $\mathcal{F}_{i}$ is defined by a logarithmic 1-form. Since each foliation $\mathcal{F}_{i}$ extends, there exists a sequence of foliations $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ in the logarithmic component $\textsf{Log}({\mathbb{P}^{N}},d_{1},d_{2},\dots,d_{k})$, where ${\left.\kern-1.2pt\mathcal{G}_{i}\vphantom{\big|}\right|_{X}}=\mathcal{F}_{i}$ and each $\mathcal{G}_{i}$ is determined by a logarithmic 1-form $\omega_{i}$. Since the integrability condition determines a proper closed set in the class space $\mathbb{P}H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{1}(k)),$ the set of points adhering to the sequence $([\omega_{i}])_{i=0}^{\infty}$ is nonempty and each point in this set is integrable. Thus, given $\omega]$ as one of these adherent points, its restriction to $X$ determines $\mathcal{F}$ by definition. Furthermore, we have that $\operatorname{codim}(\omega)\geq 2$, otherwise, $\operatorname{codim}(\operatorname{rest}(\omega))\leq 1$ and $\operatorname{rest}(\omega)$ would not determine $\mathcal{F}$. Therefore, $\omega$ determines a logarithmic foliation $\mathcal{G}$ on $\mathbb{P}^{N}$} whose restriction is $\mathcal{F}$. ∎
In general, foliations on a complete intersection $X$ extend to the projective space when the degree of the foliation is sufficiently small compared to the degrees of the hypersurfaces defining the complete intersection. In the case of hypersurfaces, [Fig23, Theorem 3.6] shows that a foliation $\mathcal{F}$ has an unique extension if $\deg(X)>2\deg(\mathcal{F})+1$. For degree-one foliations on complete intersections, we have the following result, cf. [TeseMateus, Teorema 2.1.1].
Theorem 3.2.
Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$ and dimension $n\geq 3$. If $d_{j}\geq 4$ for $j=1,\dots,s$, then any codimension-one foliation on $X$ of degree one has a unique extension to $\mathbb{P}^{n+s}$}.
Proof.
Let $\omega\in H^{0}(X,\Omega^{1}_{X}(\deg(3))$ be a section determining $\mathcal{F}$ on $X$. In light of Remark 2.1, [ACM-dist, Proposition 5.22] holds for smooth complete intersections as in the hypotheses above. Thus, the restriction map
$r_{1}\colon H^{0}({\mathbb{P}^{n+s}},\Omega^{1}_{{\mathbb{P}^{n+s}}}(3))\rightarrow H^{0}(X,\Omega^{1}_{X}(3)$
is an isomorphism, and
$r_{3}\colon H^{0}({\mathbb{P}^{n+s}},\Omega^{3}_{{\mathbb{P}^{n+s}}}(6))\rightarrow H^{0}(X,\Omega^{3}_{X}(6))$
is injective. Therefore, there exists a unique section $\alpha\in H^{0}({\mathbb{P}^{n+s}},\Omega^{1}_{{\mathbb{P}^{n+s}}}(3))$ whose restriction to $X$ is $\omega$. Note that $\alpha$ must have singularities in codimension at least two, as so does $\omega$. Observe that
$r_{3}(\alpha\wedge d\alpha)=\omega\wedge d\omega=0.$
Then $\alpha$ is integrable by the injectivity of $r_{3}$. ∎
An immediate consequence is that the space of degree-one foliations on smooth complete intersections satisfying the hypothesis of the theorem is (set-theoretically) the same as the one for the projective space, cf. [TeseMateus, Corolário 2.2.7].
Corollary 3.3.
Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$ and dimension $n\geq 3$. If $d_{j}\geq 4$ for $j=1,\dots,s$, then
${\sf Fol}(X,1)=\textsf{Log}(X;1,2)\cup\textsf{Log}(X;1,1,1).$
@@ 3.2. Codimension two foliations
For codimension-two degree-one foliations, we have the following results regarding extensions.
@@
Let $\mathcal{F}$ be a codimension-two foliation on $X$ of degree zero. In particular, $\mathcal{F}$ is a codimension-two distribution on $X$ of degree zero. Applying [TeseMateus, Corolário 2.3.3], we see that $\mathcal{F}$ has a unique extension as a codimension-two distribution to $\mathbb{P}^{n+s}$} of degree zero. Finally, by Proposition 2.4, we conclude that the extension of $\mathcal{F}$ is integrable. ∎
@@
It follows by Theorem 3.4, if $\deg(X)\geq 3$ and, for quadrics, by [AD-fano2, Proposition 3.18]. ∎
@@ 4.1. Tangential foliations
A smooth projective variety $X$ is called uniruled if, for every general point $x\in X$, there exists a non-constant morphism $f\colon{\mathbb{P}^{1}}\to X$ such that the rational curve $C=f({\mathbb{P}^{1}})$ passes through $x$ and $f^{*}{TX}$ is generated by global sections; such morphisms are called free. The scheme $\operatorname{Mor}({\mathbb{P}^{1}},X)$ parametrizing morphisms from $\mathbb{P}^{1}$} to $X$ is smooth around a free morphism $f$ and its tangent space is naturally isomorphic to $H^{0}(X,f^{*}{TX})$. We refer to [Deb-book, Kollar-book] for further details. A particular result we will need is the following.
@@
Let $X$ be a smooth uniruled projective variety, let $f\colon{\mathbb{P}^{1}}\to X$ be a free morphism, and $Z\subset X$ be a closed subvariety of codimension at least two. Then, $f^{\prime}({\mathbb{P}^{1}})\cap Z=\emptyset$ for a general deformation $f^{\prime}$ of $f$.
@@
This is a particular case of [Kollar-book, II, Proposition 3.7]. ∎
@@
See [Deb-book, Proposition 4.14] for proof of the existence of $X^{\rm free}$ over any algebraically closed field of characteristic zero. If the base field is uncountable, then $X^{\rm free}\neq\emptyset$. ∎
@@
and the natural projection $\pi:M\times{\mathbb{P}^{1}}\longrightarrow M$. We have an isomorphism $T_{M}\cong\pi_{*}\mathrm{ev}^{*}T_{X}$, and so the inclusion $T_{\mathcal{F}}\subset T_{X}$ induces a morphism $i:\pi_{*}\mathrm{ev}^{*}T_{\mathcal{F}}\hookrightarrow T_{M}$. We define the tangential foliation of $\mathcal{F}$ on $M$ by the saturation of the image of this morphism $i$ in $T_{M}$. In [LPT-def, kuster2025], the authors show how the geometry of $\mathcal{F}_{\rm tang}$ impacts the geometry of $\mathcal{F}$. We recall below two of their results.
@@
[LPT-def, Proposition 4.10] Let $X$ be a smooth uniruled projective variety and let $\mathcal{F}$ be a codimension-one foliation on $X$. If all leaves of $\mathcal{F}_{\rm tang}$ are algebraic, then there exists an algebraically integrable foliation $\mathcal{G}$ contained in $\mathcal{F}$ such that $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})=H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})$ for any sufficiently general $f\in M$. In particular, $\mathcal{F}$ is the pullback under a rational map of a foliation on a lower-dimensional manifold.
@@
[LPT-def, Theorem C] Let $X$ be a simply connected uniruled projective manifold and let $\mathcal{F}$ be a codimension-one foliation on $X$. Suppose that the general leaf of $\mathcal{F}_{\rm tang}$ is not algebraic, and a general $f]\in M$ intersects nontrivially and transversely all the algebraic hypersurfaces invariant by $\mathcal{F}$. Then $\mathcal{F}$ is defined by a closed rational $1$-form without divisorial components in its zero set.
@@
Due to Lemma 4.1, we may choose $f^{\prime}$ so that $f^{\prime}({\mathbb{P}^{1}})$ does not intersect $\operatorname{sing}(\mathcal{F})$. Aiming for a contradiction, suppose $f^{\prime}$ is tangent to $\mathcal{F}$. Recall Bott’s partial connection $\nabla^{B}\colon{T_{\mathcal{F}}}\otimes{N_{\mathcal{F}}}\to{N_{\mathcal{F}}}$ that is defined as follows. Let $v\in{T_{\mathcal{F}}}$ and $\overline{w}\in{N_{\mathcal{F}}}$ be local sections, and $w\in{TX}$ a lift of $w$. Then $\nabla^{B}_{v}(\overline{w})=\overline{[v,w]}$. Since $f^{\prime}$ is tangent to $\mathcal{F}$, we get that $f^{\prime*}\nabla^{B}$ is a holomorphic connection on the vector bundle $f^{\prime*}{N_{\mathcal{F}}}$ on $\mathbb{P}^{1}$}. By Weil’s Theorem [Atiyah-conn, Theorem 10], $f^{\prime*}{N_{\mathcal{F}}}$ must be trivial. Using that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, we conclude that $c_{1}({N_{\mathcal{F}}})=0$, hence $\mathcal{F}$ is given by a global holomorphic $p$-form $\omega\in H^{0}(X,\Omega^{p}_{X})$, $p=\operatorname{codim}\mathcal{F}$. On the other hand, since $X$ is uniruled with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, it follows that $X$ is Fano. Hence $h^{0}(X,\Omega^{p}_{X})=h^{0,p}(X)=0$, which is a contradiction. ∎
@@
Suppose $X\subset{\mathbb{P}^{N}}$ is a complete intersection given by equations of degrees $d_{1},\dots,d_{s}$. If $d_{1}+\cdots+d_{s}\leq N-1$ then $X$ is covered by lines, see [Deb-book, Proposition 2.13]. Then, the previous corollary applies.